# Jinhang Chai     Jianqing Fan

### v1

@@ 3 Low-rank-plus-sparse matrix estimation
In this section, we formulate the low-rank-plus-sparse matrix denoising problem. Consider the general observation model as follows,
@@
$\Theta^{*}$ $=L^{*}+S^{*}$ (2)
where $\Theta^{*}\in\mathbb{R}^{p\times p}$ is the parameter of interest [^4] As the work is motivated by the transition matrix estimation problem, we focus on the case that the parameter of interest is a square matrix. The extension to the general aspect ratio is relegated to the appendix. and can be written as the sum of a low-rank part $L^{*}$ and a sparse part $S^{*}$. The $\mathfrak{X}$ is a linear map. The output $Y$ can either be a vector or a matrix as long as it lives in an Euclidean space.
@@
    Matrix denoising. This corresponds to $\mathfrak{X}=\mathcal{I}$. The Markov transition matrix estimation problem is a specific instance with the parameter $\Theta^{*}$ being the ground-truth frequency matrix $F^{*}$. Letting $\widehat{F}$ be the empirical frequency matrix obtained from the data, we can define noise term as $W=\widehat{F}-F^{*}$ and write the observation model as $\widehat{F}=F^{*}+W$.
@@
    Multitask regression. Consider $p$ linear regression problems: $Y_{j}=\mathbf{X}\theta_{j}^{*}+W_{j}$ for $j\in[p]$, where $Y_{j}$ is the response vector for task $j$. Let $\Theta^{*}=(\theta_{1}^{*},\cdots,\theta_{p}^{*})$, $\mathbf{Y}=(Y_{1},\cdots,Y_{p})$, and $\mathbf{W}=(W_{1},\cdots,W_{p})$. Then, $\mathbf{Y}=\mathbf{X}\Theta^{*}+\mathbf{W}$. To reduce the number of parameters, we assume a specific low-rank-plus-sparse structure on $\Theta^{*}$ that allows us to pull the information from multiple tasks. This model corresponds to $\mathfrak{X}(\Theta)=\mathbf{X}\Theta$ in (1) with structure (2).
In order for the observation model to reveal enough information on $P^{\star}$, we make the following restricted strong convexity(RSC) condition assumption in the Frobenius norm. Following Agarwal et al., (2012), we define the regularizer $\Phi(\Delta)=\inf\limits_{L+S=\Delta}\|L\|_{*}+\lambda\|S\|_{1}$ where $\lambda$ is a given parameter, and the RSC condition as follows.
@@
From the definition of $\Phi$ we can see that $\Phi(\Delta)\leq\min(\|\Delta\|_{*},\lambda\|\Delta\|_{1})$. Therefore, Assumption 3.1 can be viewed as a strengthened variant of RSC with the restricted area being low-rank or sparse matrices.
@@
To guarantee the separability of low-rank matrices and sparse matrices, following Candes & Plan, (2010), we define incoherent matrices in Definition 3.3 and impose Assumption 3.4 on $L^{\star}$ and $S^{\star}$ in which we assume that the upper bounds are known for $\mu$, $r$ and $s$: $\bar{\mu}\geq\mu,\ \bar{r}\geq r,\ \bar{s}\geq s$.
@@
A matrix $U\in\mathcal{O}_{p,r}$ is defined to be $\mu$-incoherent if $\|U\|_{2,\infty}\leq\sqrt{\frac{\mu r}{p}}$ and denote as $U\in\mathcal{O}_{p,r}^{\mu}$. A rank-r matrix $M\in\mathbb{R}^{p\times p}$ with SVD $M=U\Sigma V^{\top}$ is defined to be $\mu$-incoherent if both $U$ and $V$ are both $\mu$-incoherent.
@@
The low-rank matrix $L^{\star}$ and sparse matrix $S^{\star}$ satisfy the following condition.
@@
Now we are ready to introduce our incoherent-constrained least-square estimator as follows:
@@ 3.1 Results on low-rank-plus-sparse matrix estimation
Let $\Delta_{L}=\widehat{L}-L^{\star},\Delta_{S}=\widehat{S}-S^{\star}$. Under Assumptions 3.1, 3.2, and 3.4, we have
$\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}\leq\frac{128}{\kappa^{2}}\left(\bar{r}\|\mathfrak{X}^{\star}(W)\|^{2}+\bar{s}\|\mathfrak{X}^{\star}(W)\|_{\max}^{2}\right).$ (4)
@@
Equation (4) holds for every possible realization of noise $W$. Another notable generality of our results is that we don’t make assumptions on the magnitude ratio of $S$ and $L$ compared to the existing literature, for instance, the pervasiveness assumption in the approximate factor model literature (Fan et al.,, 2008, 2013). When $\mathfrak{X}=\mathcal{I}$, we can choose $\kappa=1$, $\tau=0$ and Assumption 3.2 holds automatically. Furthermore, when $\bar{r}\lesssim r$ and $\bar{s}\lesssim s$, we can deduce that
@@
Unlike previous works whose bound involve either the condition number or incoherence parameter, the clean deterministic bound (5) is optimal in the following sense.
@@
$e(w):=\sup_{Y}\inf_{\widehat{L},\widehat{S}=f(Y,w)}\sup_{(L,S)\in\mathcal{C}:Y=L+S+W,\|W\|_{F}\leq w}\rho(L,S,\widehat{L},\widehat{S}),$
@@
If we can find two ways of decomposing $Y=L_{1}+S_{1}+W_{1}=L_{2}+S_{2}+W_{2}$ such that $(L_{1},S_{1}),(L_{2},S_{2})\in\mathcal{C}$ and $\|W_{1}\|_{F},\|W_{2}\|_{F}\leq w$, then we say $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})$ is a “valid confusion tuple” at the noise level $w$. As any algorithm only knowing $Y$ and $w$ is incapable of distinguishing between these two cases, we can bound the deterministic lower bound as
@@
Proposition 3.6.
For every $W$, let $w=\|W\|_{F}$. There exists a valid confusion tuple $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})$ such that $\rho(L_{1},S_{1},L_{2},S_{2})=r\|W\|^{2}$; there also exists another valid tuple $(L_{1}^{\prime},L_{2}^{\prime},S_{1}^{\prime},S_{2}^{\prime},W_{1}^{\prime},W_{2}^{\prime})$ such that $\rho(L_{1}^{\prime},S_{1}^{\prime},L_{2}^{\prime},S_{2}^{\prime})=s\|W\|_{\max}^{2}$. Therefore, we have $e(w)\geq\{r\|W\|^{2}+s\|W\|_{\max}^{2}\}/{8}$.
@@
Let $W=U\Sigma V^{T}$ be the SVD of $W$. We construct the first valid confusion tuple as $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})=(\|W\|UV^{T},0,0,0,W-\|W\|UV^{T},W)$. It is easy to verify that $(L_{1},S_{1}),(L_{2},S_{2})\in\mathcal{C}$ and $\|W_{1}\|_{F},\|W_{2}\|_{F}\leq w$, as $W_{1}$ can be viewed as a projection residue of $W_{2}=W$ to the rank-$r$ matrix space.
Similarly we construct the second valid confusion tuple as $(L_{1}^{\prime},L_{2}^{\prime},S_{1}^{\prime},S_{2}^{\prime},W_{1}^{\prime},W_{2}^{\prime})=(0,0,S_{1}^{\prime},0,W-S_{1}^{\prime},W)$, where $S_{1}^{\prime}=W\mathcal{I}(|W|\geq t)$ for some $t$, is the projection of $W$ to the sparse matrix space with sparsity level at most $s$. It is easy to verify again that $(L_{1}^{\prime},S_{1}^{\prime}),(L_{2}^{\prime},S_{2}^{\prime})\in\mathcal{C}$ and $\|W_{1}^{\prime}\|_{F},\|W_{2}^{\prime}\|_{F}\leq w$. ∎
@@
    In the case when $W$ is sampled independently from Gaussian with variance parameter $\sigma^{2}$, the minimax risk in this setting, denoted by $\mathfrak{R}^{1}$, satisfies $\mathfrak{R}^{1}\asymp\max\{p,q\}r\sigma^{2}$ if $s\leq(p+q)r/\log(p+q)$.
@@
    In the case when $W$ is supported on arbitrary predetermined $s_{0}\leq s$ entries and each entry is sampled independently from a Gaussian with variance parameter $\sigma^{2}$, let $E$ be the index matrix that takes $1$ on the predetermined support. If $s\gtrsim r\|E\|$, the minimax risk in this setting, denoted by $\mathfrak{R}^{2}$, satisfies $\mathfrak{R}^{2}\asymp s\sigma^{2}$. In particular, when $E$ is incomplete permutation matrix(each row and column has at most one $1$), then the condition on $s$ boils down to $s\gtrsim r$.
In both cases, the upper bound and lower bound match, and our estimator (3) is minimax optimal.
@@
The independence assumption on the noise can be loosened to weakly independent continuous sub-Gaussian. An example is Zhang & Wang, (2019) where the noise has the form of empirical probability error. The technique should be modified only in calculating the KL divergence which is approximately of the same order as the independent case.
@@ 3.2 A Block Alternating Minimization Algorithm
The above problem can be solved by using tools like Augmented Lagrangian method or others that are guaranteed to find a stationary point. However, in general (6) is still hard to solve due to nonconvexity, and there are no guarantees for finding global optima. Here we propose to use the alternating minimization method to solve (3) which is simpler to implement and exhibits good performance in our experiments. We slightly rewrite the original problem (3) as
@@
Note that there are four blocks of variables $U,\Sigma,V,S$ in the above formulation and each one is subject to light constraints [^5] The constraints include orthogonal normality, incoherence and sparsity.. Alternating minimization algorithm proceeds as consecutively optimizing over one block variable while keeping the others fixed.
@@
where $\mathcal{T}_{\bar{s}}$ stands for the hard thresholding operator that only keeps the $\bar{s}$ elements of the matrix with largest magnitude and zeros out other elements. Mathematically, let $|X_{\sigma(1)}|\geq|X_{\sigma(2)}|\geq\cdots,|X_{\sigma(p^{2})}|$ be the ranking of $X$’s elements in absolute value. Then  [^6] If there are multiple entries whose absolute value is equal to $|X_{\sigma(\bar{s})}|$, then we let the thresholding process keeps exactly $\bar{s}$ nonzero elements, where ties can be broken arbitrarily.
@@
For the remaining two block variables, by symmetry it suffices to consider the update formula of $U$. Specifically, $\arg\min_{U}f(U,\Sigma,V,S)$ can be formulated as
@@
One way is to view it as a standard QCQP(Quadratic Constraints Quadratic Programming) and we can solve it using standard methods. Alternatively, we can discard the incoherent constraint which therefore admits an explicit iteration scheme, viz.,
@@
The details of the algorithm is elaborated below in Algorithm 1.
@@
2:Initialize $U_{0}=\Sigma_{0}=V_{0}=S_{0}=0$.
3:for $k=1,2,\cdots,K$ do
4:     Update $S$ as $S_{k}:=\mathcal{T}_{\bar{s}}(Y-U_{k-1}\Sigma V_{k-1})$ where $\mathcal{T}_{\bar{s}}$ is given by (7).
5:     Update $\Sigma$ as $\Sigma_{k}:=\operatorname{Diag}(U_{k-1}^{\top}(Y-S_{k})V_{k-1})$.
6:     if using Method1 then
7:         Update $U$ as $U_{k}:=\operatorname{sign}((Y-S_{k})V_{k-1}\Sigma_{k})$.
8:         Update $V$ as $V_{k}:=\operatorname{sign}((Y-S_{k})^{\top}U_{k}\Sigma_{k})$.
@@
10:     if using Method2 then
11:         Update $U$ as $U_{k}=\arg\max_{U}\langle(Y-S_{k})V_{k-1}\Sigma_{k},U\rangle\quad\text{s.t.}\quad\|U\|_{2,\infty}\leq\bar{\mu},\ U^{\top}U=I$.
12:         Update $V$ as $V_{k}=\arg\max_{V}\langle(Y-S_{k})^{\top}U_{k}\Sigma_{k},V\rangle\quad\text{s.t.}\quad\|V\|_{2,\infty}\leq\bar{\mu},\ V^{\top}V=I$.
13:     end if
14:     if $\mathcal{C}$ holds then
15:         Set $U_{K}=U_{k},\Sigma_{K}=\Sigma_{k},V_{K}=V_{k},S_{K}=S_{k}$ and terminate.
16:     end if
17:end for
18:Output: $\widehat{L}=U_{K}\Sigma_{K}V_{K}^{T}$ and $\widehat{S}=S_{K}$.

### v2

@@ 2 Low-rank-plus-sparse matrix estimation
In this section, we formulate the low-rank-plus-sparse matrix denoising problem. Consider the general observation model:
@@
$\Theta^{\star}$ $=L^{\star}+S^{\star}$ (2)
where $\Theta^{\star}\in\mathbb{R}^{p\times p}$ is the parameter of interest [^4] As the work is motivated by the transition matrix estimation problem, we focus on the case that the parameter of interest is a square matrix. The extension to the general aspect ratio is relegated to the supplementary material. and can be written as the sum of a low-rank part $L^{\star}$ and a sparse part $S^{\star}$. The $\mathfrak{X}$ is a linear map. The output $Y$ can either be a vector or a matrix as long as it lives in an Euclidean space.
@@
    Matrix denoising. This corresponds to $\mathfrak{X}=\mathcal{I}$. The Markov transition matrix estimation problem is a specific instance with the parameter $\Theta^{\star}$ being the ground-truth frequency matrix $F^{\star}$. Letting $\widehat{F}$ be the empirical frequency matrix obtained from the data, we can define noise term as $W=\widehat{F}-F^{\star}$ and write the observation model as $\widehat{F}=F^{\star}+W$.
@@
    Multitask regression. Consider $p$ linear regression problems: $Y_{j}=\mathbf{X}\theta_{j}^{\star}+W_{j}$ for $j\in[p]$, where $Y_{j}$ is the response vector for task $j$. Let $\Theta^{\star}=(\theta_{1}^{\star},\cdots,\theta_{p}^{\star})$, $\mathbf{Y}=(Y_{1},\cdots,Y_{p})$, and $\mathbf{W}=(W_{1},\cdots,W_{p})$. Then, $\mathbf{Y}=\mathbf{X}\Theta^{\star}+\mathbf{W}$. To reduce the number of parameters, we assume a low-rank-plus-sparse structure on $\Theta^{\star}$ that allows us to pull the information from multiple tasks. This model corresponds to $\mathfrak{X}(\Theta)=\mathbf{X}\Theta$ in (1) with structure (2).
In order for the observation model to reveal enough information on $\Theta^{\star}$, we make the following restricted strong convexity (RSC) condition assumption in the Frobenius norm. Following the same notation of Agarwal et al., (2012), we define the regularizer $\Phi(\Delta)=\inf\limits_{L+S=\Delta}\|L\|_{*}+\lambda\|S\|_{1}$ where $\lambda$ is a given parameter, and the RSC condition is as follows.
@@
The RSC establishes a form of approximate identifiability. From the definition of $\Phi$ we can see that $\Phi(\Delta)\leq\min(\|\Delta\|_{*},\lambda\|\Delta\|_{1})$. Therefore, Assumption 2.1 can be viewed as a strengthened variant of RSC with the restricted area being low-rank or sparse matrices. In later proofs, we make use of the fact that the nuclear norm and 1-norm are the dual norm of 2-norm and max-norm, respectively.
@@
To guarantee the separability of low-rank matrices and sparse matrices, following Candes & Plan, (2010), we define incoherent matrices in Definition 2.3 and impose Assumption 2.4 on $L^{\star}$ and $S^{\star}$ in which we assume that the upper bounds are known for $\mu$, $r$ and $s$: $\bar{\mu}\geq\mu,\ \bar{r}\geq r,\ \bar{s}\geq s$. The incoherence condition ensures that the singular vectors of a low-rank matrix are not overly concentrated in any particular directions or entries, which is crucial for matrix completion (Candes & Recht,, 2012; Chen,, 2015). In our setting, it also facilitates separability from the sparse matrix.
@@
A matrix $U\in\mathcal{O}_{p,r}$ is defined to be $\mu$-incoherent if $\|U\|_{2,\infty}\leq\sqrt{\frac{\mu r}{p}}$ and denote as $U\in\mathcal{O}_{p,r}^{\mu}$. A rank-r matrix $M\in\mathbb{R}^{p\times p}$ with SVD $M=U\Sigma V^{\top}$ is defined to be $\mu$-incoherent if both $U$ and $V$ are $\mu$-incoherent.
@@
There exists some constant $c>0$ such that the low-rank matrix $L^{\star}$ and sparse matrix $S^{\star}$ satisfy the following conditions:
@@
Now we introduce our incoherent-constrained least-square estimator as follows:
@@ 2.1 Results on low-rank-plus-sparse matrix estimation
Let $\Delta_{L}=\widehat{L}-L^{\star},\Delta_{S}=\widehat{S}-S^{\star}$. Under Assumptions 2.1, 2.2, and 2.4, we have
$\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}\leq\frac{128}{\kappa^{2}}\left(\bar{r}\|\mathfrak{X}^{*}(W)\|^{2}+\bar{s}\|\mathfrak{X}^{*}(W)\|_{\max}^{2}\right).$ (4)
@@
Inequality (4) holds for every possible realization of noise $W$. Another notable generality of our results is that we don’t make assumptions on the magnitude ratio of $S$ and $L$ compared to the existing literature, for instance, the pervasiveness assumption in the approximate factor model literature (Fan et al.,, 2008, 2013). When $\mathfrak{X}=\mathcal{I}$, we can choose $\kappa=1$, $\tau=0$ and Assumption 2.2 holds automatically. Furthermore, when $\bar{r}\lesssim r$ and $\bar{s}\lesssim s$, we can deduce that
@@
The proof of Theorem 2.5 relies significantly on the following key separation lemma, with its detailed proof provided in the supplementary material.
Lemma 2.6.
Let $P=U\Sigma V^{\top}$ and $Q=L\Lambda R^{\top}$ be the SVD of $P,Q\in\mathbb{R}^{p\times p}$, $U,V,L,R\in\mathcal{O}_{p,r}$. Denote $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ and $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}\geq 0$ as the diagonal elements of $\Sigma$ and $\Lambda$, respectively. Suppose $P$ and $Q$ both satisfy the $\mu$-incoherence condition 2.3, then we have for $\Delta=P-Q$,
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\leq\frac{\tilde{c}\mu r^{3}}{p}$ (6)
where $\tilde{c}$ is a fixed constant independent of both $P$ and $Q$.
A detailed comparison with existing works under similar settings is given in Appendix A in the supplementary material. For example, see Table 1 there. Unlike previous works whose bound involves either the condition number or incoherence parameter, the clean deterministic bound (5) is optimal in the following sense.
@@
$e(w):=\inf_{\widehat{L},\widehat{S}=f(Y,w)}\sup_{(L,S)\in\mathcal{C},\|W\|_{F}\leq w}\rho(L,S,\widehat{L},\widehat{S}),$
@@
If we can find two ways of decomposing $Y=L_{1}+S_{1}+W_{1}=L_{2}+S_{2}+W_{2}$ such that $(L_{1},S_{1}),(L_{2},S_{2})\in\mathcal{C}$ and $\|W_{1}\|_{F},\|W_{2}\|_{F}\leq w$, then we say $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})$ is a “valid confusion tuple” at the noise level $w$. As any algorithm only knowing $Y$ and $w$ is incapable of distinguishing between these two cases, we can bound the deterministic lower bound. More rigorously, for fixed $Y,w$, define
$\widehat{L},\widehat{S}=\arg\inf_{\widehat{L},\widehat{S}=f(Y,w)}\sup_{(L,S)\in\mathcal{C}:Y=L+S+W,\|W\|_{F}\leq w}\rho(L,S,\widehat{L},\widehat{S}),$
then we have
@@
Theorem 2.7.
For every $W$, let $w=\|W\|_{F}$. There exists a valid confusion tuple $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})$ such that $\rho(L_{1},S_{1},L_{2},S_{2})=r\|W\|^{2}$; there also exists another valid tuple $(L_{1}^{\prime},L_{2}^{\prime},S_{1}^{\prime},S_{2}^{\prime},W_{1}^{\prime},W_{2}^{\prime})$ such that $\rho(L_{1}^{\prime},S_{1}^{\prime},L_{2}^{\prime},S_{2}^{\prime})=s\|W\|_{\max}^{2}$. Therefore, we have
$e(w)\geq\{r\|W\|^{2}+s\|W\|_{\max}^{2}\}/{8}.$
@@
Let $u,v$ be the left and right singular vectors corresponding to the singular value of $\|W\|$. We construct the first valid confusion tuple as $(L_{1},L_{2},S_{1},S_{2},W_{1},W_{2})=(\|W\|uv^{T},0,0,0,W-\|W\|uv^{T},W)$. It is easy to verify that $(L_{1},S_{1}),(L_{2},S_{2})\in\mathcal{C}$ and $\|W_{1}\|_{F},\|W_{2}\|_{F}\leq w$, as $W_{1}$ can be viewed as a projection residue of $W_{2}=W$ to the rank-$r$ matrix space.
Similarly, we construct the second valid confusion tuple. Define ${\cal I}(|W|\geq t)\in\mathbb{R}^{p\times q}$ with ${\cal I}(|W_{ij}|\geq 0)$ as the $(i,j)$-element. Similarly define $W{\cal I}(|W|\geq t)\in\mathbb{R}^{p\times q}$ with $W_{i,j}{\cal I}(|W_{ij}|\geq 0)$ as the $(i,j)$-element. Let $(L_{1}^{\prime},L_{2}^{\prime},S_{1}^{\prime},S_{2}^{\prime},W_{1}^{\prime},W_{2}^{\prime})=(0,0,S_{1}^{\prime},0,W-S_{1}^{\prime},W)$, where $S_{1}^{\prime}=W\mathcal{I}(|W|\geq t)$ for some $t$, is the projection of $W$ to the sparse matrix space with sparsity level at most $s$. It is easy to verify again that $(L_{1}^{\prime},S_{1}^{\prime}),(L_{2}^{\prime},S_{2}^{\prime})\in\mathcal{C}$ and $\|W_{1}^{\prime}\|_{F},\|W_{2}^{\prime}\|_{F}\leq w$. ∎
@@
The minimax risks for two noise distributions are given as follows.
@@
    Consider the case when each row of $W$ is sampled independently from multivariate Gaussian: $W_{i,:}\sim{\cal N}(0,\Sigma_{i})$ with $c\sigma^{2}I\leq\Sigma\leq C\sigma^{2}I$. Then the minimax risk in this setting, denoted by $\mathfrak{R}^{1}$, satisfies $\mathfrak{R}^{1}\asymp\max\{p,q\}r\sigma^{2}$ if $s\leq(p+q)r/\log(p+q)$, where we recall that $p\times q$ are the dimensions of $W$.
@@
    Consider the case when $W$ is supported on arbitrary predetermined $s_{0}\leq s$ entries, and each entry is sampled independently from a Gaussian with variance parameter $\sigma^{2}$. Let $E$ be the index matrix that takes $1$ on the predetermined support. If $\|E\|\lesssim\sqrt{\frac{s}{r}}$, the minimax risk in this setting, denoted by $\mathfrak{R}^{2}$, satisfies $\mathfrak{R}^{2}\asymp s\sigma^{2}$, up to logarithmic factors. In particular, when $E$ is an incomplete permutation matrix(each row and column has at most one $1$), then the condition on $s$ boils down to $s\gtrsim r$.
In both cases, the upper bound and lower bound match, and estimator (3) is minimax optimal.
@@
Note that the first case includes a special subcase where all entries of $W$ are independent Gaussian ${\cal N}(0,\sigma^{2})$. The independence assumption on the noise can be loosened to weakly independent continuous sub-Gaussian. An example is Zhang & Wang, (2019) where the noise has the form of empirical probability error. The technique should be modified only in calculating the KL divergence which is approximately of the same order as the independent case. Also, for the noise matrix with independent rows but dependent across columns (Agterberg et al.,, 2022), as long as the covariance matrices have upper and lower bounded eigenvalues, the upper bound is retained.
Remark 5.
For simplicity, we take $\mathfrak{X}$ to be identity. It can be readily extended to $\mathfrak{X}$ satisfying Assumption 2.1 and that $\|\mathfrak{X}\|_{op}\leq C_{op}$. The proof of changes in bounding upper bound by $\|\mathfrak{X}^{*}\|_{op}\leq C_{op}$ and calculating the KL divergence of constructed instances and is omitted.
@@ 2.2 A Block Alternating Minimization Algorithm
The above problem can be solved by using tools like the Augmented Lagrangian method or others that are guaranteed to find a stationary point. However, in general (7) is still hard to solve due to nonconvexity, and there are no guarantees for finding global optima. Here, we propose to use the alternating minimization method to solve (3), which is simpler to implement and exhibits good performance in our experiments. We slightly rewrite the original problem (3) as
@@
Note that there are four blocks of variables $U,\Sigma,V,S$ in the above formulation, and each one is subject to light constraints [^5] The constraints include orthogonal normality, incoherence, and sparsity.. The alternating minimization algorithm proceeds as consecutively optimizing over one block variable while keeping the others fixed.
@@
where $\mathcal{T}_{\bar{s}}$ stands for the hard thresholding operator that only keeps the $\bar{s}$ elements of the matrix with largest magnitude and zeros out other elements. Mathematically, let $|X_{\sigma(1)}|\geq|X_{\sigma(2)}|\geq\cdots,|X_{\sigma(p^{2})}|$ be the ranking of $X$’s elements in absolute value. Then  [^6] If there are multiple entries whose absolute value is equal to $|X_{\sigma(\bar{s})}|$, then we let the thresholding process keep exactly $\bar{s}$ nonzero elements, where ties can be broken arbitrarily.
@@
For the remaining two block variables, by symmetry, it suffices to consider the update formula of $U$. Specifically, $\arg\min_{U}f(U,\Sigma,V,S)$ can be formulated as
@@
One way is to view it as a standard QCQP(Quadratic Constraints Quadratic Programming), and we can solve it using standard methods. Alternatively, we can discard the incoherent constraint, which therefore admits an explicit iteration scheme, viz.,
@@
The details of the algorithm are elaborated below in Algorithm 1.
@@
2:Initialize $U_{0}=\Sigma_{0}=V_{0}=0$.
3:if using Init1 then
4:     Initialize $S_{0}=0$.
5:else if using Init2 then
6:     Initialize $S_{0}=\mathcal{T}_{\bar{s}}(Y)$ where $\mathcal{T}_{\bar{s}}$ is given by (8).
7:else if using Random Init then
8:     Initialize $S_{0}=M_{\bar{s}}\circ Y$, where $M_{\bar{s}}$ is the mask matrix with $\bar{s}$ randomly selected elements set to $1$ and other elements set to 00.
@@
10:for $k=1,2,\cdots,K$ do
11:     if using Method1 then
12:         Update $U$ as $U_{k}:=\operatorname{sign}((Y-S_{k-1})V_{k-1}\Sigma_{k-1})$.
13:         Update $V$ as $V_{k}:=\operatorname{sign}((Y-S_{k-1})^{\top}U_{k}\Sigma_{k-1})$.
14:     else if using Method2 then
15:         Update $U$ as $U_{k}=\arg\max_{U}\langle(Y-S_{k})V_{k-1}\Sigma_{k-1},U\rangle\quad\text{s.t.}\quad\|U\|_{2,\infty}\leq\bar{\mu},\ U^{\top}U=I$.
16:         Update $V$ as $V_{k}=\arg\max_{V}\langle(Y-S_{k-1})^{\top}U_{k}\Sigma_{k-1},V\rangle\quad\text{s.t.}\quad\|V\|_{2,\infty}\leq\bar{\mu},\ V^{\top}V=I$.
17:     end if
18:     Update $\Sigma$ as $\Sigma_{k}:=\operatorname{Diag}(U_{k}^{\top}(Y-S_{k-1})V_{k})$.
19:     Update $S$ as $S_{k}:=\mathcal{T}_{\bar{s}}(Y-U_{k}\Sigma_{k}V_{k}^{\top})$.
20:     if $\mathcal{C}$ holds then
21:         Set $U_{K}=U_{k},\Sigma_{K}=\Sigma_{k},V_{K}=V_{k},S_{K}=S_{k}$ and terminate.
22:     end if
23:end for
24:Output: $\widehat{L}=U_{K}\Sigma_{K}V_{K}^{T}$ and $\widehat{S}=S_{K}$.
@@
To initialize, we typically begin with one of two methods: $S=0$ (Init1), which corresponds to starting with incomplete SVD on $Y$, or $U=\Sigma=V=0$ (Init2), which is equivalent to setting $S=\mathcal{T}_{\bar{s}}(Y)$. Alternatively, initialization can be performed randomly. Specifically, we set $S=M_{\bar{s}}\circ Y$, where $M_{\bar{s}}$ is the mask matrix with $\bar{s}$ randomly selected elements set to $1$ and other elements set to 00. These initialization schemes are included in Algorithm 1 and will be compared in Figure 1 in Section 5.
We remark in passing that while the alternating minimization algorithm (specifically, Method1) is not guaranteed to find a global maximum, it is always non-increasing, leading to the identification of a local minimum. Random initializations can be utilized, as well as warm start variations, to heuristically get rid of local minima. One can also utilize the Gumbel approximation trick applied in solving the combinatorial optimization (Gu et al.,, 2024). We will discuss this further in Section 5.
Our theory relies on knowledge of $\bar{r}$, which serves as an upper bound for $r$. For specific problems, domain knowledge can guide an informed estimate of $\bar{r}$. Alternatively, one could construct a grid of possible values for $\bar{r}$ and solve the optimization problem for each. If $\bar{r}$ is set below the true $r$, this would result in a large bias and objective function value. A comprehensive solution for selecting $\bar{r}$ is deferred to future work.